OpenStudy (anonymous):

NEED HELP suppose a parabola f(x)has its vertex at (0,1)and its zero atx=-2and x=2 then f(x)equals A.1-1/4x^2 B.4-x^2 C.x^2-1 D.x^2-4

OpenStudy (anonymous):

@Nurali can u help me ?

OpenStudy (aum):

If the two roots are 'p' and 'q', the equation of the parabola is: y = a(x-p)(x-q) where a is a constant that can be determined by substituting the vertex.

OpenStudy (anonymous):

ok can u help me solve it ?

OpenStudy (aum):

Here the two roots are 2 and -2. That means (x-2) and (x - -2) are factors. That is (x-2) and (x+2) are factors. Multiply (x-2) and (x+2). What do you get?

OpenStudy (anonymous):

96

OpenStudy (anonymous):

i plugged it in my calculator thats what i got

OpenStudy (aum):

Multiply the factors (x-2)(x+2). Use the identity: (a+b)(a-b) = a^2 - b^2 or use FOIL method to multiply it out.

OpenStudy (anonymous):

i got D as the answer is that correct

OpenStudy (aum):

Can you show the work that got you that answer?

OpenStudy (anonymous):

(x-2)(x+2)=96 ,,,,,,,,, x^2-4=96

OpenStudy (aum):

That 96 is wrong. Just multiply (x-2)(x+2) first.

OpenStudy (anonymous):

i got 96 again

OpenStudy (anonymous):

what did u get

OpenStudy (aum):

Don't use your calculator. multiplying (x-2)(x+2) will not be a number. It will be a quadratic function.

OpenStudy (anonymous):

x^2+2x+-2+4

OpenStudy (aum):

Actually it is: x^2 + 2x -2x - 4 Simplify it.